I need help with this. I am having dificulties with this. i appreciate any help I can get.

Replace the blanks in the equation with constants to complete the square and form a true equatoin.

x^2 + 2/3x+ = (x + )^2

Answer by ankor@dixie-net.com(22740)

Replace the blanks in the equation with constants to complete the square and form a true equation.
x^2 + 2/3x + ____ = (x + ___)^2
:
To complete the square, divide the coefficient of x by 2, and square it
1/2 * 2/3 = 2/6, (2/6)^2 = 4/36, so we have:
:
x^2 + 2/3x + 4/36 = (x + 2/6)^2
:
You can reduce the fractions
x^2 + 2/3x + 1/9 = (x + 1/3)^2
:
:
you can prove this to yourself
FOIL (x + 1/3)*(x + 1/3)